The countries of Utopia and Dystopia can each produce two goods: wine and cars. If Utopia produces only wine, it can produce 80 barrels per year; if it produces only cars, it can produce 40 cars per year. If Dystopia produces only wine, it can produce 60 barrels per year; if it produces only cars, it can produce 90 cars per year.

Suppose the two countries decided to trade with each other. If each country wants to take maximum advantage of the potential gains from trade, how many cars and barrels of wine should each produce?
Answer #1

To answer this question we need to first see if the countries have any comparative advantage or not. The comparative advantage is defined as the advantage in terms of opportunity cost of producing one good in terms of another. If one country can produce a good at lower opportunity cost than the other country the former country is said to have a comparative advantage.

Let's calculate the opportunity cost of producing car in both the countries in terms of wine. That is to say, how many barrels of wine needs to be sacrificed to produce 1 unit of car. That's the opportunity cost of car.

Opportunity cost of producing car in Utopia, 80/40= 2 barrels/car. That is utopia needs to sacrifice 2 barrels of wine to produce 1 car.

Opportunity cost of producing car in Dystopia, 60/90= 2/3 barrels/car. That is dystopia needs to sacrifice 2/3 barrels of wine to produce 1 car.

As we can see that opportunity cost of producing car is lower in dystopia.

Now let's calculate the opportunity cost of producing wine in each country. That how many cars are needed to be sacrificed in order to produce 1 barrel of wine.

Opportunity cost of producing wine in utopia, 40/80= 1/2 cars/barrel. That utopia needs to sacrifice 1/2 a unit of car to produce 1 barrel of wine.

Opportunity cost of producing wine in dystopia, 90/60= 3/2cars/barrel, That is dystopia needs to sacrifice 3/2 units of car to produce 1 barrel of wine.

As we can see that utopia has lower opportunity cost of producing wine.

Now to take maximum advantage of potential gains from trade a country must specialize in producing a good in which it has a comparative advantage in terms of Lower opportunity cost. If each country produces the good in which it has a comparative advantage the global output will be maximized and gains from trade will be maximized too.

So, to maximize the potential gains from trade utopia must specialize in producing wine and dystopia must specialize in producing cars.

Feedback will be appreciated.